# Higher order Jacobi method for solving a system of linear equations

### v1

@@ 4 Discussions
The computational complexity of solving matrix equations depends primarily on two factors [35, 36, 37], the sparsity of the coefficient matrices and the type of method used, direct or iterative. The system matrices built using PDEs are usually sparse with very few nonzero elements in each row and are diagonally dominant. The computational complexity of solving such sparse systems is much lower when compared with their denser counterparts, such as Hamiltonian matrices in computational quantum chemistry, due to the reduced number of calculations that can be performed in each row. A comprehensive comparison of different computational complexities is shown in Table 2.
@@
The proposed higher-order method has the same computational complexity as matrix-matrix multiplication. Even though it is an order of magnitude higher than the matrix-vector multiplication, unlike iterative operations, the matrix-matrix multiplication is a highly parallelizable operation [44, 45, 46, 47].
Moreover, the theoretical computational complexity of multiplying large matrices has improved significantly over the years, falling from default $O(n^{3})$ to the current known limit of $O(n^{2.376})$ [42] for large matrices.
From Figs. 1, 2, and 3, it is clear that the actual computation times of the higher-order method while generating the higher-order coefficients and even solving matrix equations using the generated coefficients are higher than the built-in matrix equation solving methods of the software packages used. However, the order of complexity of generating coefficients is better than the built-in matrix solve methods of any package, which suggests an algorithmic improvement in solving matrix equations. This also suggests that the proposed higher-order method is highly parallelized, making it exceptionally well suited for leveraging advancements in parallel computing.
As coefficient matrices become denser with recursive squaring, pruning techniques [48, 49] used in traditional neural networks can be adopted to make dense higher-order matrices much more compact. However, with recent developments in parallel computing and the ever-increasing affordability of GPUs [50, 51] and even photonic circuits [52], matrix-matrix multiplications can be accelerated.
The proposed higher-order method for matrix inversion structurally aligns with a deep linear network. Once the coefficients are explicitly derived for a known system, the method enables efficient resolution of variations of that system without the need for computationally intensive re-computation of the coefficients. The recurrent iterative scheme inherently adopts the form of a deep recurrent neural network. This framework aims to provide analytical insight into the explicit definition of network parameters, thereby circumventing the training process and significantly reducing computational overhead.
@@ 5 Conclusion
Physical systems are fundamentally governed by differential equations, and recent advances have shown growing interest in Physics-Informed Neural Networks (PINNs) for solving such equations. However, in most existing frameworks, the term “physics-informed” pertains primarily to the incorporation of physical laws into the training objective, rather than to the explicit definition of network parameters such as weights, biases, and activation functions.
This work presents a direction toward interpreting and defining the inner workings of neural networks through explicit analytical formulation. In particular, a modified Jacobi iteration method is developed to compute higher-order coefficient matrices using matrix-matrix multiplications. The resulting formulation, referred to as the higher-order Jacobi method, exhibits the structural characteristics of a shallow linear network with a limited number of layers. Through this formulation, it is also possible to compute the inverse of the original coefficient matrix directly via matrix operations, eliminating the need for iterative inversion procedures.
In addition, an iterative solution scheme is proposed whereby the network parameters, once determined for a known system, can be reused to solve variations of the system without re-computing the coefficients. This capability is especially advantageous in applications where the underlying system undergoes parametric changes. The iterative process takes the form of a deep recurrent neural network, offering a principled and efficient approach that maintains consistency across related systems.
Computational analysis indicates that the proposed methods lead to a measurable improvement in the order of complexity. This signifies a potential algorithmic advancement in solving matrix equations more efficiently, with reduced computational cost and improved scalability. The presented methodology not only connects traditional numerical methods with neural architectures but also contributes toward interpretable [53, 54] and training-free network designs grounded in physical and mathematical principles.
Future directions include the extension of this framework to non-linear systems and time-dependent differential equations, enabling broader applications in scientific computing and model-driven learning.

### v2

@@ 4 Discussions
The computational complexity of solving matrix equations depends primarily on two factors [35, 36, 37], the sparsity of the coefficient matrices and the type of method used, direct or iterative. The system matrices built using PDEs are usually sparse with very few nonzero elements in each row and are diagonally dominant. The computational complexity of solving such sparse systems is much lower when compared with their denser counterparts, such as Hamiltonian matrices in computational quantum chemistry. This is due to the reduced number of calculations performed in each row. A comprehensive comparison of different computational complexities is shown in Table 3.
@@
The proposed higher-order method has the same computational complexity as matrix-matrix multiplication. Although its naive approach is an order of magnitude higher than matrix-vector multiplication, matrix-matrix multiplication is a highly parallelizable operation [45, 46, 47, 48].
Moreover, the theoretical computational complexity of multiplying large matrices has improved significantly over the years, falling from the default $O(n^{3})$ to the current known limit of $O(n^{2.376})$ [43] for large matrices. However, practical and scalable implementations even with the help of deep learning algorithms [42, 49]still tend to perform more closely to the Strassen algorithm in practice.
However, the higher order Jacobi method takes advantage of the parallelizability of matrix-matrix multiplication, unlike the current sequential approaches, and therefore demonstrates exceptional performance when implemented on GPUs.
As coefficient matrices become denser with recursive squaring, pruning techniques [50, 51] used in traditional neural networks can be adopted to make dense higher-order matrices much more compact. However, with recent developments in parallel computing and the ever-increasing affordability of GPUs [52, 53] and even photonic circuits [54], matrix-matrix multiplications can be further accelerated.
The proposed higher-order method for matrix inversion structurally aligns with a deep linear network. Once the coefficients are explicitly derived for a known system, the method enables efficient resolution of variations of that system without the need for computationally intensive re-computation of the coefficients. The recurrent iterative scheme inherently adopts the form of a deep recurrent neural network. This framework aims to provide analytical insight into the explicit definition of network parameters in physics informed neural networks, thereby circumventing the training process and significantly reducing computational overhead.
@@ 5 Conclusion
Physical systems are fundamentally governed by differential equations, and recent advances have shown growing interest in Physics-Informed Neural Networks (PINNs) for solving such equations. However, in most existing frameworks, the term “physics-informed” relates primarily to the incorporation of physical laws into the training objective, rather than to the explicit definition of network parameters such as weights, biases, and activation functions.
This work presents a direction toward interpreting and defining the inner workings of neural networks through an explicit analytical formulation. In particular, a modified Jacobi iteration method is developed to compute higher-order coefficient matrices using matrix-matrix multiplications. The resulting formulation, referred to as the higher order Jacobi method, exhibits the structural characteristics of a shallow linear network with a limited number of layers. Through this formulation, it is also possible to compute the inverse of the original coefficient matrix directly via matrix operations, eliminating the need for iterative inversion procedures.
In addition, an iterative scheme is proposed in which the network parameters, once determined for a known system, can be reused to solve variations of the system without recomputing the coefficients. This capability is especially advantageous in applications where the underlying system undergoes parametric changes. The iterative process takes the form of a deep recurrent neural network, offering a principled and efficient approach that maintains consistency across related systems.
Computational analysis indicates that the proposed methods lead to exceptional improvement in the order of complexity when implemented on GPU. This signifies a significant algorithmic advancement in solving matrix equations, with reduced computational cost and improved scalability. The presented methodology not only connects traditional numerical methods with neural architectures, but also contributes to interpretable [55, 56] and training-free network designs grounded in physical and mathematical principles.
Future directions include the extension of this framework to nonlinear systems and time-dependent differential equations, and eigenvalue problems, enabling broader applications in scientific computing and model-driven learning.
Code Availability
A simple implementation of the proposed method and the code used to generate the data and the plots are available at the following link: https://github.com/nithingoona/HOJM